OpenStudy (anonymous):

Simplify the expression. 12p5 A. 11,880 B. 95,040 C. 1,235,520 D. 7,920

OpenStudy (skullpatrol):

What is "p"?

OpenStudy (anonymous):

I have no idea.. ^p

OpenStudy (skullpatrol):

12p5?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):

Maybe $$12p^5$$

OpenStudy (anonymous):



OpenStudy (anonymous):

yes I believe

jimthompson5910 (jim_thompson5910):

n P r = (n!)/((n-r)!) 12 P 5 = (12!)/((12-5)!) 12 P 5 = ???

OpenStudy (anonymous):

divide it?

jimthompson5910 (jim_thompson5910):

what is 12! equal to

OpenStudy (anonymous):

to 5?

OpenStudy (anonymous):

12P5=12!/(12-5)! = 95040

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

95040 is correct, nice work

jimthompson5910 (jim_thompson5910):

the steps leading up to 95040 are also correct

OpenStudy (anonymous):

Thanks :)